# Neural Concept Verifier: Scaling Prover-Verifier Games via Concept Encodings

### v2

@@ Appendix A Experimental Details: Baselines
In this section we provide training details for the three introduced baselines ResNet-18, Pixel-MAC and CBM.
A.1 ResNet-18
@@
For this setup, we initialize the framework with a pretrained ResNet-18 model and employ the Adam optimizer across all experiments. For the CLEVR datasets, we utilize a batch size of 128 and train for 30 epochs with a learning rate of $10^{-4}$ and weight decay of $10^{-4}$. The model is trained using 20 different random seeds, with early stopping based on the validation loss.
@@
For CIFAR-100, we train the ResNet-18 model for 300 epochs using a learning rate of $10^{-4}$ and weight decay of $10^{-4}$, with a batch size of 128. The model is trained across 10 random seeds using early stopping based on validation loss.
@@
For ImageNet, the same setup is applied with a lower learning rate of $10^{-5}$, while keeping the batch size, weight decay, and early stopping criterion unchanged. Both datasets use the standard pixel-level inputs and all trained checkpoints are reused to initialize subsequent Pixel-MAC experiments.
A.2 Pixel-MAC
@@
For ImageNet, we further reduce the batch size to 32 due to memory constraints and train for 80 epochs using the same learning rates and hyperparameters as in the CIFAR-100 setting. The mask size is set to 1000 pixels per image. As with CIFAR-100, early stopping is applied across 10 random seeds, and training is initialized from the pretrained ResNet-18 backbone.
A.3 CBM with Linear Classifier
@@
For CIFAR-100 and ImageNet, we train a linear classifier on sparse SpLiCE encodings using a dictionary size of 10,000. In both cases, we use a batch size of 4096 and train for 250 epochs with early stopping, a learning rate of $10^{-3}$, and no weight decay. A hidden layer with 512 units are used, and an L1 penalty of 0.2 is applied within SpLiCE to encourage sparsity in the concept representations. All results are averaged over 10 random seeds.
Appendix B Experiments: NCB-based Neural Concept Verifier
@@
B.1 Pretraining
Before training NCV, we first pretrain the models without the feature selectors. The corresponding results for the pretraining are shown in Table  3, where we evaluate on 20 random seeds. These pretrained models are then used as initialization for the subsequent NCV training. For the pretraining, we use a Set Transformer with two stacked multi-head attention blocks, a hidden dimension of 128 and four attention heads. We use a batchsize of 128, 30 epochs and the Adam optimizer with a learning rate of $10^{-3}$ for both datasets, applying early stopping based on validation loss.
@@ B.1 Pretraining
B.2 NCV Training
For the experiments presented in our main results in Table  1, the experimental details for both datasets are as follows:
@@ B.2 NCV Training
Extended Results. Additionally, we evaluated the NCV framework with varying mask sizes and an alternative model architecture for the feature selectors. The results are presented in Table  4 for the CLEVR-Hans3 data set and Table  5 for the CLEVR-Hans7 dataset, where we evaluate both the validation set and the test set. The alternative architecture implements a MLP with two hidden layers and ReLU activation functions for the feature selectors, while maintaining a pretrained Set Transformer as the classifier across all experiments. Our results reveal that the Set Transformer feature selector consistently outperforms the MLP feature selector on test set, particularly with smaller mask sizes such as 4 and 6. Furthermore, this configuration maintains high completeness (>96%) and soundness (>99%), even with a reduced number of selected features.
@@
MLP $97.41\pm 1.07$ $99.99\pm 0.03$ $97.01\pm 0.93$ $99.99\pm 0.04$
B.3 Explanations
Here, we present supplementary examples of explanations generated by both Pixel-MAC and NCV on the CLEVR-Hans3 dataset in Figure  4. These results further substantiate our claim that NCV provides significantly more transparent and interpretable explanations compared to the pixel-based PVG baseline.
@@ B.3 Explanations
B.4 Shortcut Mitigation
In Table  6, we provide complementary results to those in Table  2, evaluating model robustness on CLEVR-Hans3 under varying degrees of confounding. As in the main experiment, models are trained on datasets with increasing ratios of non-confounded samples, and evaluated on both a confounded validation set and a non-confounded test set.
We observe similar trends: NCV consistently achieves the highest test accuracy across all settings, and its validation–test gap narrows more rapidly than for both the linear and non-linear CBM baselines. Even with minimal clean data (e.g., 5%), NCV achieves strong generalization and a markedly smaller gap. These findings reinforce our conclusions from CLEVR-Hans7 and further highlight NCV’s ability to mitigate shortcut learning through its concept-level PVG formulation.
Table 6: NCV shows increased robustness to shortcuts on CLEVR-Hans3. We report validation accuracy on a set with, and test accuracy on a set without shortcuts, as well as the validation–test gap (lower is better) across models trained with varying amounts of clean data. A smaller gap indicates improved generalization and reduced reliance on shortcuts. NCV consistently achieves the highest test accuracy and among the lowest gaps, confirming its ability to learn robust, faithful decision boundaries under the presence of shortcuts.
Ratio Clean (Samples) Model Val Acc (w/ shortcut) Test Acc (w/o shortcut) Val-Test Gap (↓)
[…]
NCV 99.37± $\mathbf{99.37}\mbox{\scriptsize$\pm\,0.18$}$ 97.11± $\mathbf{97.11}\mbox{\scriptsize$\pm\,0.98$}$ $\mathbf{2.26}$
[…]
NCV 99.59± $\mathbf{99.59}\mbox{\scriptsize$\pm\,0.19$}$ 98.88± $\mathbf{98.88}\mbox{\scriptsize$\pm\,0.37$}$ $\mathbf{0.71}$
[…]
NCV 99.37± $\mathbf{99.37}\mbox{\scriptsize$\pm\,0.28$}$ 98.82± $\mathbf{98.82}\mbox{\scriptsize$\pm\,0.67$}$ $\mathbf{0.55}$
Appendix C Experimental Details: CLIP-based Neural Concept Verifier
@@
C.1 Pretraining
Once more, before starting with the actual NCV training, we first pretrain the models without the provers (Merlin and Morgana). The corresponding results for the pretraining are shown in Table  7. As textual concept descriptions $T$, we used the top 10,000 most frequent one- and two-word phrases from LAION (Schuhmann et al.,, 2021) captions, following the setup and codebase of Bhalla et al., (2024). For pretraining the verifier, we use a two-layer multilayer perceptron (MLP) with a hidden dimension of 512 and GELU activations (Hendrycks and Gimpel,, 2016) on both CIFAR-100 and ImageNet-1k. For CIFAR-100 and ImageNet, we train the verifier using a batch size of 4096 and a learning rate of $10^{-4}$, applying dropout with a rate of 0.3 and a weight decay of $10^{-4}$. Models are trained for 300 epochs with early stopping (patience 10) based on validation loss. All pretraining is conducted without provers, and the resulting verifiers are used to initialize the CLIP-based NCV training.
@@ C.1 Pretraining
Table 7: Pretraining accuracy of the verifier (without provers) for CLIP-based NCV on CIFAR-100 and ImageNet-1k.
C.2 NCV Training
For the experiments presented in our main results in Table  1, we detail the training setup separately for CIFAR-100 and ImageNet.
@@ C.2 NCV Training
CIFAR-100. For CIFAR-100, all components are trained jointly for 100 epochs using the Adam optimizer. We set the verifier learning rate to $10^{-4}$, and use $5\times 10^{-4}$ for both Merlin and Morgana. A batch size of 256 is used throughout. Weight decay is set to $0.1$, and a hard mask size of 32 concepts is enforced per input. To incentivise sparse masks, an L1 penalty of 0.1 is applied to the provers. A learning rate scheduler (plateau-based) is employed with a patience of 5, minimum learning rate of $10^{-6}$, and decay factor of 0.001. Early stopping is disabled, and all results are averaged over 10 random seeds.
@@
[…]
@@
[…]

### v3

@@ Appendix C Experimental Details: Baselines
In this section, we provide training details for the considered baselines: ResNet-18, ResNet-50, Pixel-MAC, CBM (linear and nonlinear variants), and DCR.
C.1 ResNet-18 and ResNet-50
@@
We initialize the framework with a pretrained ResNet-18 model and employ the Adam optimizer across all experiments. On the CLEVR datasets, the model is trained with a batch size of 128 for 30 epochs using a learning rate of $10^{-4}$ and weight decay of $10^{-4}$, repeated across 20 random seeds with early stopping based on validation loss.
@@
On CIFAR-100, we use a ResNet-50 trained for 100 epochs with a learning rate of $10^{-4}$, weight decay of $10^{-5}$, and a batch size of 128, averaged over 10 random seeds with early stopping. The ResNet-50 baseline on ImageNet is evaluated directly using pretrained PyTorch (Paszke et al., 2019) weights without further finetuning. For COCOLogic, a ResNet-50 is trained for 300 epochs with a batch size of 256, learning rate of $10^{-4}$, and weight decay of $10^{-2}$, again averaged over 10 random seeds with early stopping.
@@
In the Pixel-MAC setup, a separate ResNet-18 is trained under the same configuration as above but with a reduced learning rate of $10^{-5}$, while keeping the batch size, weight decay, and early stopping criterion unchanged. All Pixel-MAC results are obtained from these ResNet-18 checkpoints.
C.2 Pixel-MAC
@@
For ImageNet and COCOLogic, we further reduce the batch size to 32 due to memory constraints and train for 80 epochs using the same learning rates and hyperparameters as in the CIFAR-100 setting. The mask size is set to 1000 pixels per image. As with CIFAR-100, early stopping is applied across 10 random seeds, and training is initialized from the pretrained ResNet-18 backbone.
C.3 CBM with Linear Classifier
@@
For CIFAR-100 and ImageNet, we train a linear classifier on sparse SpLiCE encodings using a dictionary size of 10,000. In both cases, we use a batch size of 4096 and train for 250 epochs with early stopping, a learning rate of $10^{-3}$, and no weight decay. A hidden layer with 512 units is used, and an L1 penalty of 0.2 is applied within SpLiCE to encourage sparsity in the concept representations. All results are averaged over 10 random seeds.
C.4 DCR Baseline Implementation
For the additional nonlinear CBM comparison in Sec.  4, we implemented Deep Concept Reasoning (DCR; Barbiero et al. 2023) following the official torch-explain library. Our implementation uses the same experimental setup as NCV to ensure a fair comparison.
Architecture.
We employ a ResNet-18 backbone (pretrained weights disabled for consistency) followed by two DCR-specific modules: (1) a ConceptEmbedding layer that maps backbone features to concept predictions $\mathbf{c}_{\text{pred}}\in[0,1]^{C}$ and learned concept embeddings $\mathbf{c}_{\text{emb}}\in\mathbb{R}^{C\times d}$, and (2) a ConceptReasoningLayer that learns class-specific logic rules in Disjunctive Normal Form (DNF) over the concept embeddings to produce class predictions.
Concept Supervision.
We supervise the concept predictions using the same CLIP-Sim similarity vectors as NCV. Since these are cosine similarities in [1,1][-1,1], we apply a linear normalization $c_{\text{truth}}=(\text{sim}+1)/2$ to map them to $[0,1]$ for BCE loss compatibility.
Training.
Following the official DCR tutorial, we optimize a joint loss $\mathcal{L}=\mathcal{L}_{\text{concept}}+\lambda\cdot\mathcal{L}_{\text{task}}$, where both terms use binary cross-entropy (BCE). The task loss uses one-hot encoded labels, as the ConceptReasoningLayer outputs per-class probabilities (not logits). We train with Adam optimizer for 100 epochs.
Hyperparameter Sweep.
For each dataset, we performed a grid search over:
  • Vocabulary size: $n\in\{1000,3000,10000\}$
  • Concept embedding dimension: $d\in\{8,16,32\}$
  • Learning rate: $\eta\in\{10^{-3},10^{-4}\}$
  • Task loss weight: $\lambda\in\{0.5,1.0\}$
This resulted in $3\times 3\times 2\times 2=36$ configurations per dataset, totaling over 100 training runs. For COCOLogic-10, we report balanced accuracy due to class imbalance.
Results.
Tab.  4 summarizes the best accuracies achieved by DCR compared to NCV at 10k vocabulary size. DCR successfully trains on CIFAR-100 and COCOLogic-10 but fails to complete training on ImageNet-1k within reasonable time (>3>3 days per epoch). In contrast, NCV’s sparsity-inducing mechanism enables efficient scaling to all three datasets under the same 10k-concept space.
Table 4: Comparison between DCR and NCV at 10k-concept vocabulary size on CLIP-Sim. NCV metrics correspond to completeness (accuracy). For COCOLogic-10, we report balanced accuracy due to class imbalance.
Dataset DCR NCV (ours)
(acc. %) (completeness, %)
CIFAR-100 51.76 83.32
ImageNet-1k n/a (timeout) 67.04
COCOLogic-10 (bal.) 38.61 75.42
Appendix D NCB-based Neural Concept Verifier Experiments
@@
D.1 Pretraining
Before training NCV, we first pretrain the models without the feature selectors. The corresponding results for the pretraining are shown in Tab.  5, where we evaluate on 20 random seeds. These pretrained models are then used as initialization for the subsequent NCV training. For the pretraining, we use a Set Transformer with two stacked multi-head attention blocks, a hidden dimension of 128 and four attention heads. We use a batch size of 128, 30 epochs and the Adam optimizer with a learning rate of $10^{-3}$ for both datasets, applying early stopping based on validation loss.
@@ D.1 Pretraining
D.2 NCV Training
For the experiments presented in our main results in Tab.  1, the experimental details for both datasets are as follows:
@@ D.2 NCV Training
Extended Results. Additionally, we evaluated the NCV framework with varying mask sizes and an alternative model architecture for the feature selectors. The results are presented in Tab.  6 for the CLEVR-Hans3 dataset and Tab.  7 for the CLEVR-Hans7 dataset, where we evaluate both the validation set and the test set. The alternative architecture implements a MLP with two hidden layers and ReLU activation functions for the feature selectors, while maintaining a pretrained Set Transformer as the classifier across all experiments. Our results reveal that the Set Transformer feature selector consistently outperforms the MLP feature selector on the test set, particularly with smaller mask sizes such as 4 and 6. Furthermore, this configuration maintains high completeness (>96%) and soundness (>99%), even with a reduced number of selected features.
@@
MLP $7.41\pm 1.07 99$$9.99\pm 0.03$ $7.01\pm 0.93 99$$9.99\pm 0.04$
D.3 Explanations
Here, we present supplementary examples of explanations generated by both Pixel-MAC and NCV on the CLEVR-Hans3 dataset in Fig.  4. These results further substantiate our claim that NCV provides significantly more transparent and interpretable explanations compared to the pixel-based PVG baseline.
@@ D.3 Explanations
Appendix E Experimental details for CLIP-based NCV
@@
E.1 Pretraining
Once more, before starting with the actual NCV training, we first pretrain the models without the provers (Merlin and Morgana). The corresponding results for the pretraining are shown in Tab.  8. As textual concept descriptions $T$, we used the top 10,000 most frequent one- and two-word phrases from LAION (Schuhmann et al., 2021) captions, following the setup of Bhalla et al. (2024). For pretraining the verifier, we use a two-layer multilayer perceptron (MLP) with a hidden dimension of 512 and GELU activations (Hendrycks and Gimpel, 2016) on CIFAR-100, ImageNet-1k, and COCOLogic. On CIFAR-100 and ImageNet, training uses a batch size of 4096 and a learning rate of $10^{-4}$, with dropout (0.3), weight decay of $10^{-4}$, and early stopping (patience 10). On COCOLogic, we instead train for 100 epochs with a batch size of 512, learning rate of $10^{-4}$, and weight decay of $10^{-2}$, using a learning-rate scheduler (plateau, patience 5, factor $10^{-3}$, minimum learning rate $10^{-6}$) and no early stopping. All pretraining is conducted without provers, and the resulting verifiers are used to initialize the CLIP-based NCV training.
Table 8: Pretraining accuracy of the verifier (without provers) for CLIP-based NCV on CIFAR-100, COCOLogic and ImageNet-1k.
@@ E.1 Pretraining
E.2 NCV Training
For the experiments presented in our main results in Tab.  1, we detail the training setup separately for CIFAR-100 and ImageNet.
@@ E.2 NCV Training
CIFAR-100. For CIFAR-100, all components are trained jointly for 100 epochs using the Adam optimizer. We set the verifier learning rate to $10^{-4}$, and use $5\times 10^{-4}$ for both Merlin and Morgana. A batch size of 256 is used throughout. Weight decay is set to $0.1$, and a hard mask size of 32 concepts is enforced per input. To incentivize sparse masks, an L1 penalty of 0.1 is applied to the provers. A learning rate scheduler (plateau-based) is employed with a patience of 5, minimum learning rate of $10^{-6}$, and decay factor of 0.001. Early stopping is disabled, and all results are averaged over 10 random seeds.
@@
COCOLogic. Training on COCOLogic follows the CIFAR-100 configuration with minor adjustments: models are trained for 100 epochs with a batch size of 512, verifier learning rate of $10^{-4}$, and prover learning rates of $5\times 10^{-4}$. We use a weight decay of $0.01$, a mask size of 32, and an L1 penalty of 0.1 on the provers. As with CIFAR-100, early stopping is disabled, and learning rate scheduling and seed averaging remain unchanged.
@@
16 $94.08\mbox{\scriptsize$\pm\,1.46$}$ $81.82\mbox{\scriptsize$\pm\,0.50$}$ $9.96\mbox{\scriptsize$\pm\,0.03$} 99$$9.97\mbox{\scriptsize$\pm\,0.02$}$
@@
4 $9.30\mbox{\scriptsize$\pm\,0.34$} 55$$5.96\mbox{\scriptsize$\pm\,0.30$}$ $9.81\mbox{\scriptsize$\pm\,0.03$} 99$$9.83\mbox{\scriptsize$\pm\,0.06$}$
@@
E.3 Prover and Verifier Architectures
In all experiments, the choice of prover and verifier architectures is guided by the structure of the concept representation. For unordered, slot-based concept encodings (as in NCB on CLEVR-Hans), we use permutation-invariant Set Transformers for Merlin, Morgana, and Arthur. For fixed-size concept vectors (as in CLIP/SpLiCE on CIFAR-100, ImageNet-1k, and COCOLogic-10), we use shallow nonlinear MLPs.
In preliminary experiments, we observed that performance is stable across a range of nonlinear architectures (varying depth, width, and activations), whereas purely linear models consistently underperformed by collapsing to simple correlation tests. Based on this, we recommend using permutation-invariant architectures for slot-based concepts and nonlinear MLPs for vector-based concepts when applying NCV to new datasets.
E.4 Effect of the weighting parameter $\gamma$
The weighting parameter $\gamma$ controls the trade-off between Merlin’s cooperative objective and Morgana’s adversarial objective when training Arthur. Recall that Arthur’s loss is given by
$L_{A}=(1-\gamma)\,L_{M}+\gamma\,L_{\widehat{M}}$,
so that $\gamma=0$ corresponds to training Arthur only on Merlin’s loss, and larger $\gamma$ increases the relative weight of Morgana’s adversarial objective. To assess the influence of this trade-off on NCV, we sweep $\gamma$ over a range from 0 (Merlin-only) up to $0.5$ (substantial weight on Morgana) on CIFAR-100, ImageNet-1k, and COCOLogic, keeping all other hyperparameters fixed. For each setting, we train three models with different random seeds and report the mean completeness and soundness; for COCOLogic, we report balanced metrics due to class imbalance.
Across all datasets the same qualitative behavior emerges. When $\gamma=0$, i.e., the verifier is trained without adversarial pressure from Morgana, completeness remains high but soundness collapses (e.g., $37.9\%$ on CIFAR-100, $10.4\%$ on ImageNet-1k, and $51.8\%$ balanced soundness on COCOLogic). As soon as $\gamma>0$, soundness rapidly recovers to values close to those reported in Tab.  1, while completeness remains essentially unchanged throughout the range of $\gamma$ we consider. These results demonstrate that incorporating adversarial concept selection is crucial for learning verifiers that remain reliable under misleading concept subsets, without sacrificing standard predictive performance in the regimes we study. For a detailed discussion of why an adversarial prover and its relative strength are essential for the Merlin–Arthur mutual-information guarantees underlying NCV, see Suppl.  A.
(a) CIFAR-100
(b) ImageNet-1k (validation)
(c) COCOLogic (validation, balanced metrics)
Figure 5: Effect of the weighting parameter $\gamma$ on completeness and soundness for (a) CIFAR-100, (b) ImageNet-1k validation, and (c) COCOLogic validation (balanced metrics). All curves show means over 3 random seeds.
Appendix F Computational Cost and Hardware Setup
To complement the main results, we report approximate training times for all models and datasets considered in Sec.  4. All runs were executed on a single NVIDIA A100 or A40 GPU; times are reported as rounded wall-clock estimates and are meant to convey relative cost rather than exact benchmarks. For CBM baselines, we exclude the per-sample SpLiCE optimization step at inference time, which would add substantial overhead and further increase their deployment cost.
Table 11: Approximate training time comparison across models and datasets. Times are rounded wall-clock estimates on a single NVIDIA A100 or A40 GPU. For CBM baselines, the per-sample SpLiCE optimization at inference time is omitted here, but would incur significant additional cost. (ImgNet = ImageNet-1k, COCO-10 = COCOLogic-10, CL-H3/CL-H7 = CLEVR-Hans3/7.)
Model CIFAR ImgNet COCO-10 CL-H3 CL-H7
ResNet (baseline) $\sim$25–30m $\sim$1.5d $\sim$50m <10<10m <10<10m
Pixel-MAC $\sim$1.3d $\sim$3d $\sim$4h <2<2h <2<2h
CBM $\sim$10–15m $\sim$4h $\sim$2m $\sim$2m $\sim$2m
NCV (ours) $\sim$20m  (11m A + 7m PVG) $\sim$5h  (2.5h A + 2.5h PVG) $\sim$5m $\sim$1.5m $\sim$5m
Overall, NCV is 1133 orders of magnitude cheaper to train than Pixel-MAC across all datasets, while remaining comparable to standard CBMs in runtime. Since concept encodings are precomputed once and reused across models, NCV scales similarly to a conventional classifier without additional architectural overhead, and remains practical even in large-scale settings such as ImageNet-1k.
Appendix G Use of Large Language Models
Large language models were used to support this work by assisting with text refinement, implementation of code components (including methods and plot generation), and by providing input during idea development and approach refinement.